Summary statistics for returns on two stocks X and Y are listed below.

Mean Variance
Stock X 5.81% 0.004000
Stock Y 3.11% 0.007000

The covariance of returns on stocks X and Y is 0.002100. Consider a portfolio of 10% stock X and 90% stock Y.

What is the variance of portfolio returns?

Please round your answer to six decimal places.

Note that the correct answer will be evaluated based on the full-precision result you would obtain using Excel.
Answer #1
Variance = (wxơx)2 + (wyơy)2 + 2*Cov(x,y)wxwy =0.1^2*0.004+0.9^2*0.007+2*0.1*0.9*0.0021= 0.006088